OpenStudy (anonymous):

a gardener wants to make a rectangular enclosure using a wall as one side and 160m of fencing for the other three sides, as shown. For what values of x is the area of the enclosure more than 3000m^2 but less than 3150m^2?

OpenStudy (anonymous):



OpenStudy (anonymous):

STEP 1: MAKE A DIAGRAM done STEP 2: MAKE EQUATIONS 2x+l=160 x*l=area STEP 3: SOLVE FOR X (or l) x*l=3000 2x+l=160 ...